OpenStudy (anonymous):

sin4L=4sinLcosL(1-2sin^2L)

OpenStudy (anonymous):

use the cosine and sin double angle formulas cos(2l)=cos^2(l)-sin^2(l)=1-2sin^2(l) sin(2l)=2sinlcosl

OpenStudy (anonymous):

sin(20)=sin(10+10)=2sin(10)cos(10)

OpenStudy (anonymous):